OpenStudy (anonymous):

Find the general solution of tan 3x = 3 tanx

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

tan 3x = ? ... expand this, solve the cubic or quartic equation for tan

OpenStudy (anonymous):

Let me try by that way

OpenStudy (anonymous):

\[\tan 3x = \frac{ 3tanx - \tan^3 x }{ 1-3\tan^2x }\]

OpenStudy (anonymous):

\[\frac{ 3\tan - \tan^3 x }{ 1-3 \tan^2 x } = 3tanx\] \[\tan^3 x = 9\tan^3 x\]

OpenStudy (experimentx):

wolf says solution is n(pi) http://www.wolframalpha.com/input/?i=tan%283x%29+%3D+3+tan%28x%29

OpenStudy (anonymous):

But hw...???

OpenStudy (experimentx):

must be quartic in tan(x)

OpenStudy (anonymous):

Nope...!

hartnn (hartnn):

8 tan^3 x= 0 tan x = 0

OpenStudy (experimentx):

tan(3x) = should have cubic ... you have only square ( quadratic)

hartnn (hartnn):

tan x=0 only at n*pi

OpenStudy (experimentx):